For the year 20X1, the Landmark Restaurant had sales of $800,000 and expenses of $700,000 excluding depreciation. The building is being leased and the only depreciable asset is equipment in the amount of $100,000. The equipment has a life of ten years with zero salvage value.

Required:

Calculate the earnings before taxes for 20X1 if:

1. The equipment is depreciated under straight-line depreciation

2. The equipment is depreciated under double-declining balance depreciation
Answer #1

Answer

  • Requirement 1

>Working

A

Cost

$          100,000.00

B

Residual Value

$                           -  

C=A - B

Depreciable base

$          100,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$            10,000.00


>Earnings before taxes = $ 800000 sales - $ 700000 expense - $ 10000 depreciation = $ 90,000

  • Requirement 2

>Working

A

Cost

$          100,000.00

B

Residual Value

$                           -  

C=A - B

Depreciable base

$          100,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$            10,000.00

F=E/C

SLM Rate

10.00%

G=F x 2

DDB Rate

20.00%

  • Earnings before taxes = $ 800000 sales - $ 700000 expense – ($100000 x 20%) depreciation
    = 800000 – 700000 – 20000
    = $ 80,000